The number of positive integral solutions of XYZT=210 are


A)256


b)64


c)24


d)120


this is the problem given in free iitjee 2010 practice papers in askiitians website


and in the solution u din't include 1 as the factor which is also a positive integer.


please reply me on this


 

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago
Dear siri given XYZT =210 factor of 210 is 2*3*5*7 so total number of solution = 4*4*4*4 = 256 this is the answer i got for this question Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best. Regards, Askiitians Experts Badiuddin